Von Bora Corporation is expected to pay a dividend of $1.40 per share at the end of this year and a $1.50 per share at the end of the second year. You expect Von Bora's stock price to be $25.00 at the end of two years. Von Bora's equity cost of capital is 10%. Suppose you plan on purchasing Von Bora stock in one year, right after the $1.40 dividend is paid. You then plan on selling your stock at the end of year two, right after the $1.50 dividend is paid.

The capital gain rate that you will receive on your investment is closest to:

The dividend yield that you will receive on your investment is closest to:
Answer #1

Purchase price = present value of the future cash flows (Dividend + Price in year 2)

Purchase price = (1.5 + 25)/1.1 = 24.0909090909

Capital Gain = 25/24.0909090909 - 1 = 0.03773584906

Capital Gain = 3.773584906%

Dividend yield = 1.5/ 24.0909090909 = 6.226415094%

Can you please upvote? Thank You :-)